The velocity of a particle executing SHM is 0.12 ms^-1and 0.04 ms^-1 when its displacements from the mean position are 0.01m and 0.03m respectively. Calculate the amplitude, period and maximum velocity of the particle ?
1
Expert's answer
2019-09-30T10:21:26-0400
"v_1=\omega \sqrt{a^2-x_1^2},v_2=\omega \sqrt{a^2-x_2^2}"

"\frac{v_1}{v_2}=\frac{\sqrt{a^2-x_1^2}}{\sqrt{a^2-x_2^2}}"

"\frac{0.12}{0.04}=\frac{\sqrt{a^2-0.01^2}}{\sqrt{a^2-0.03^2}}"

"a=0.032\ m"

"v_1=\omega \sqrt{a^2-x_1^2}\to 0.12=\omega \sqrt{0.032^2-0.01^2}"

"\omega=4\frac{rad}{s}"

"v_{max}=\omega a=(4)(0.032)=0.13\frac{m}{s}"

"T=\frac{2\pi}{\omega}=1.6\ s"